A girl pulls a 10kg wooden cart with a constant force of 30N. What is the cart acceleration?


1
Expert's answer
2021-04-19T17:06:29-0400

According to 2nd Newton's law, the acceleration is "a = \frac{F}{m} = \frac{30 N}{10 kg} = 3 \frac{m}{s^2}".